Grade 12th passPhysical Chemistry

50 ml 0.1 M NaOH is added to 50 ml of 0.1 M CH3COOH solution, what will be the final pH? please give me a detailed solution for this. Given ka=1.8×10^-5

Pawan joshi
7 Years agoGrade 12th pass

1 Answer

Arun
7 Years ago
after adding 50ml 0.1M NaOH in 50ml of 0.1M CH3COOH solution,there willbe made 50ml 0.1 M CH3COONa & there remaining 50ml 0.1M CH3COOH. total velocity=(50+50+50)=150ml.Molarity of CH3COOH=(50x0.1)/150=.0333M.Molarity of CH3COONa=(50x0.1)/150=.0333M.then pH=pKa+log [salt]/[acid]=4.7447+log [.0333]/[.0333]=4.7447+log 1=4.7447+0=4.7447.So the ans is pH=4.7447.
 
Regards
Arun